OpenStudy (anonymous):

A rectangle has an area represented by x^2-4x-12 square feet. If the length is x+2 feet, what is the width of the rectangle?

OpenStudy (anonymous):

The area of a rectangle is defined as the product of its length and width, \(A=lw\). It should make sense that we can factor our area polynomial into the length and width, just as you can factor a rectangle with area five into its dimensions 5x1. To factor \(A=x^2-4x-12\) we find factors of -12 which sum to -4... i.e. -6 and 2. Our linear factors must turn to 0 for these values of \(x\). we can factor A into \((x+6)(-2)\). If \(x-2\) is the length then the width must be \(x+6\).

OpenStudy (anonymous):

(x+6)(x-2)... sorry, I'm on an iPad.

OpenStudy (anonymous):

Omg thanks ;-;

OpenStudy (anonymous):

But wouldn't that be the other way around since 2 times -6 is -4? x+2 and x-6

OpenStudy (anonymous):

No; to make \((x-2)(x+6)=0\), we need to make either \(x-2=0\) or \(x+6=0\). Solve both for \(x\) to yield \(x=2\) and \(x=-6\).